OpenStudy (anonymous):

Factor the expression into a product of binomials. 25a^2-70ab+49b^2 Is this the same as before??

OpenStudy (anonymous):

I have 4 different choices

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

is it going to be (25a+49b)(25a-49b)?

OpenStudy (2bornot2b):

\[25a^2-70ab+49b^2\implies (5a-7b)^2\implies(5a-7b)(5a-7b)\]

OpenStudy (2bornot2b):

thats wrong sinz

OpenStudy (anonymous):

how do we get the 70 what = it

OpenStudy (anonymous):

\( (5 a-7 b)^2 \)

OpenStudy (anonymous):

5 * 7 = 35, since you have two of these, double it to get 70

OpenStudy (anonymous):

okay that's what I was wondering is if you doubled the 35?

OpenStudy (2bornot2b):

\[25a^2-70ab+49b^2\implies25a^2-(2\times5a\times 7b)+49b^2\implies(5a-7b)^2\implies(5a-7b)(5a-7b)\]

OpenStudy (anonymous):

your equation ran off the screen just to let you know

OpenStudy (2bornot2b):

It comes from the famous formula...\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

^it's not a formula, just expansion ;)

OpenStudy (2bornot2b):

what exactly is the difference between the two terms?

OpenStudy (anonymous):

so does this go with the same equation then 25x^2- 60xy+36y^2

OpenStudy (2bornot2b):

Not arguing though.. I would love to know

OpenStudy (anonymous):

haha, it's okay, IMO formula is abstract term, where as expansion is much more concrete ;)

OpenStudy (2bornot2b):

@sinz yes the expression \[25x^2- 60xy+36y^2\] can be dealt in a similar way

OpenStudy (2bornot2b):

@fool for math, can you show me an example so that I understand better.

OpenStudy (anonymous):

okay so it would be (5x - 6y)^2 right?

OpenStudy (2bornot2b):

@sinz yes you are right

OpenStudy (anonymous):

Formula is something that you will like to memorize, expansion is more like understanding what's happening pertaining to the above binomial expansion.

OpenStudy (anonymous):

SOme people memorize it as a formula and then understand the expansion

OpenStudy (anonymous):

others, understand the expansion and then try to generalize it as a formula.

OpenStudy (anonymous):

btw the binomial expansion or formula whatever you call it, it's originated from the study of the expansion and then trying to somehow form a closed a form

OpenStudy (anonymous):

you can call the closed form as formula, but unless you understand the expansion it's not worth-awhile memorizing the formula.

OpenStudy (2bornot2b):

Okay okay, got the hint. Thanks @foolformath

OpenStudy (anonymous):

This was how pascal triangles was formed .. and further research on identities like hockey sticks and depth analysis of combinatorics started .

OpenStudy (anonymous):

it's fantastic actually, now if you just memorize it as a formula you are missing a whole lot of mathematics :D

OpenStudy (anonymous):

Glad to help :)